OpenStudy (my4babiesandme):

how do i solve r=124sqrt8000/10

OpenStudy (anonymous):

technically it is already solved. all you can do from here is simplify. is that what you want done?

OpenStudy (my4babiesandme):

During the first part of a trip, a canoeist travels 80 miles at a certain speed. The canoeist travels 19 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 5 hrs. What was the speed on each part of the trip? round to nearest hundredth. this is the problem i am doing and that is the step in i can understand the example to

OpenStudy (dumbcow):

Equations \[80 = r*t_{1}\] \[19 = (r-5)t_{2}\] \[t_{1} +t_{2} = 5\] \[\rightarrow \frac{80}{r} + \frac{19}{r-5} = 5\] \[80(r-5) + 19r = 5r(r-5)\] \[5r^{2} -124r+400 = 0\] \[r = \frac{124 \pm \sqrt{7376}}{10}\]

OpenStudy (dumbcow):

From here just use a calculator, ignore answer where r < 5 \[r = 20.99\]

OpenStudy (my4babiesandme):

ok what i do not understand is how to go from the last step what do i do to get the 20.99?

OpenStudy (dumbcow):

use a calculator \[\sqrt{7376} = 85.8836\] \[124 + 85.8836 = 209.8836\] \[\frac{209.8836}{10} = 20.99\]

OpenStudy (my4babiesandme):

thank you so much you are awesome

OpenStudy (dumbcow):

no problem

OpenStudy (my4babiesandme):

i so understand it now i have done 3 problems to make sure thank you so much